# LLM-Based Detection of Tangled Code Changes for Higher-Quality Method-Level Bug Datasets

### v2

@@ 4.3. RQ3: Performance of embedding-based machine learning models
This research question explores the viability of embedding-based ML models for detecting tangled code changes. Unlike prompt-based approaches that rely on in-context reasoning of LLMs, this method uses embeddings, vector representations of the input data from commit messages and code diffs, to train supervised classifiers. The key objective is to assess whether these embeddings, when used as features in ML models, can achieve competitive performance in detecting bug-related code changes. This experiment is motivated by the previous studies where embedding-based models performed well and even outperformed prompt-based techniques (Petukhova et al., 2024; Keraghel et al., 2024; Kokkodis et al., 2025).
@@
Figure 3. The methodology for RQ3 has two main parts: (i) embedding generations and (ii) building ML models.
[…]
Overview of methodology for embedding generation.
To generate embeddings, we utilized Gemini’s text embedding model, text-embedding-004, with the Classification task type, which provides strong performance and adaptability across NLP tasks, particularly in generating embeddings optimized for classification scenarios (Lee et al., 2024). This model yields 768-dimensional vector representations for each input instance, constructed by concatenating the code diff and its commit message. Due to constraints in resources, this study focuses exclusively on this proprietary LLM. To reduce cost, gpt-4o and gpt-4o-mini were excluded, and Gemini’s free quota was used to generate the embeddings. We acknowledge that we may achieve higher accuracy using GPT models. Nonetheless, as the objective is to assess the general effectiveness of embedding-based models, the results offer a representative indication of the potential performance of similar LLM embeddings.
The classification task was treated as a binary problem, and the dataset was split into an 80% training set and a 20% test set. We evaluated multiple classifiers, including RandomForest (RF), Support Vector Classifier (SVC), DecisionTree (DT), and Multi-Layer Perceptron (MLP). Table 4 presents the performance comparison of all models, showing that the MLPClassifier achieved the best results, with an accuracy of 0.8952, F1 score of 0.9019, and MCC of 0.7911, outperforming traditional classifiers in all metrics.
@@
Table 4. Performance comparison of ML models trained on LLM embeddings. Metrics are calculated by targeting the buggy samples (pos_label = ”Buggy”).
Classifier Accuracy Precision Recall F1 Score MCC
RF 0.8159 0.8413 0.8196 0.8303 0.6294
SVC 0.8839 0.9180 0.8660 0.8912 0.7684
DT 0.6856 0.7173 0.7062 0.7117 0.3660
MLP 0.8952 0.9290 0.8763 0.9019 0.7911
@@
Furthermore, we evaluated the MLPClassifier using a Leave-One-Out (LOO) strategy to test its robustness, where a single example is used as the test sample and the remaining serve as the training data. This approach not only offers more robust evaluation, but also can show if a model’s accuracy can be improved with more training samples. Encouragingly, the model achieved improved performance with an accuracy of 0.9036, precision of 0.8980, recall of 0.9141, F1 Score of 0.9060, and MCC of 0.8073, setting a new benchmark compared to prior best results.
In terms of class-wise performance (Table 5), the model maintained a balanced detection capability, with precision and recall exceeding 0.89 for both classes, further validating its discriminative strength across both Buggy and NotBuggy samples.
Table 5. Class-wise performance of MLPClassifier.
Class Precision Recall F1 Score
Buggy 0.90 0.91 0.91
NotBuggy 0.91 0.89 0.90
These results indicate that embedding-based ML models, particularly deep neural networks like MLP, are highly effective for the task of tangled change detection. Their performance not only surpasses traditional rule-based or heuristic methods but also rivals that of LLM-based prompt engineering strategies.
@@
Summary of RQ3: Embedding-based ML models, particularly MLPClassifier, demonstrate strong predictive power for tangled change detection, achieving over 90% F1 score and outperforming prior baselines. More encouragingly, our results suggest that the performance of embedding-based models can be improved with an enlarged training dataset.
@@ 4.4. RQ4: Potential impact of LLM-based untangling on future method-level bug prediction models
From the results of RQ1 to RQ3, we observe that the LLM-based method-level tangled change detection approach performs well with high accuracy and F1 scores using both prompting techniques and embedding-based ML models. This suggests that such approaches can effectively reduce noise in bug datasets. However, an important question remains: Does this noise reduction benefit machine learning-based bug prediction models? Building such a model, however, requires addressing several challenges. First, a broad set of code metrics is needed, whereas the dataset from Chowdhury et al. (Chowdhury et al., 2024) includes only five. Second, Mashhadi et al. (Mashhadi et al., 2023) demonstrate that combining code metrics with embeddings leads to better performance, requiring us to generate embeddings for all methods. Finally, a thorough evaluation across multiple algorithms and fine-tuning strategies is essential. Due to these requirements, constructing a full bug prediction model is beyond the scope of this paper. Nonetheless, if we can demonstrate that our noise reduction approach improves the ability of code metrics to distinguish between bug-prone and non-bug-prone methods, likely, the performance of machine learning models relying on these metrics will also improve. Therefore, this research question examines whether the distribution of code metrics becomes more distinguishable in the dataset produced by our method.
In RQ1 and RQ2, we evaluated multiple prompting strategies and LLMs, showing that GPT-4o combined with few-shot + chain-of-thought prompting delivers the best performance. In RQ3, we examined the feasibility of using LLM-generated embeddings with downstream machine learning classifiers to detect tangled changes. Although the latter approach yields better results, it requires generating embeddings for all tangled code diffs from bug-fix commits to support prediction with the machine learning model, which is more complex and time-consuming than the few-shot + chain-of-thought approach. Therefore, we employed GPT-4o with few-shot + chain-of-thought prompting to detect tangled changes across all methods in 49 software projects in a fast and efficient manner.
In the dataset by Chowdhury et al. (Chowdhury et al., 2024), each method is associated with its change history and the number of other methods modified in the same commit. We apply our few-shot and chain-of-thought prompting strategy using GPT-4o to separate NotBuggy changes within bug-fix commits, thereby removing noise. To enable replication, the algorithm used to generate the Less-Noisy dataset is presented in Algorithm 1. After constructing the Less-Noisy dataset, we obtain four types of methods:
@@
Algorithm 1 Creation of Less-Noisy Dataset by Untangling.
1:for all project files do
2:    for all methods do
3:         if method age $<2$ years then
4:             continue
5:         end if
6:         if any commit is buggy (High-Prec.) and tangled then
7:             for all commits do
8:                 if buggy and tangled then
9:                     detect with LLM
10:                     if detection is “Buggy” then
11:                         mark Buggy, break
12:                     end if
13:                 else if buggy and not tangled then
14:                     mark Buggy, break
15:                 end if
16:             end for
17:             if none marked Buggy then
18:                 mark NotBuggy
19:             end if
20:         else if any commit is buggy (High-Prec.) then
21:             mark Buggy
22:         else if any commit is not buggy (High-Recall) then
23:             mark NotBuggy
24:         end if
25:    end for
26:end for
The dataset provided by Chowdhury et al. (Chowdhury et al., 2024) contains five code metrics computed for each version of a method across its history. For our analysis, we consider the metrics from the first version of each method to ensure consistency across the dataset. These code metrics are defined as follows: Size refers to the number of source lines of code, excluding comments and blank lines (Chowdhury et al., 2022, 2024); Readability is a score that reflects the ease of reading the source code, developed by Buse et al. (Buse and Weimer, 2009); McCabe measures cyclomatic complexity by counting the number of independent execution paths in a method (McCabe, 1976); FanOut represents the number of distinct methods called by a given method, indicating its dependency footprint (Pascarella et al., 2020; Chowdhury et al., 2024); and Maintainability Index (MI) is a composite metric that combines several aspects, including complexity and size, to offer a comprehensive assessment of code maintainability (Oman and Hagemeister, 1992).
@@
Figure 4. Distribution of code metrics for Buggy and NotBuggy methods in the Titan project across the Noisy and Less-Noisy datasets. Metric distributions differ more significantly in the Less-Noisy dataset.
[…]
Boxplot of Titan
@@
To further quantify the differences between Buggy and NotBuggy methods, we compute Cliff’s Delta effect sizes across all 49 projects. Table 6 presents the effect size distributions for both the Noisy and Less-Noisy datasets. We categorize the effect sizes following Hess et al. (Hess and Kromrey, 2004): Negligible ($<0.147$), Small ($0.330.147\leq\delta<$ 0.33), Medium ($0.4740.33\leq\delta<$ 0.474), and Large ($\delta\geq 0.474$).
@@
The table reports the proportion of projects falling into each effect size category for all five code metrics. The results indicate that, for most metrics, the Less-Noisy dataset shows a substantial shift toward larger effect sizes. For the code metric Size, in the Less-Noisy dataset, 73% of the projects exhibit a large effect size in the difference between Buggy and NotBuggy methods, compared to only 36% in the Noisy dataset. Similar patterns are observed for the other four code metrics. These findings clearly demonstrate that our filtering approach effectively eliminates noise and increases the distinction between the code metrics of buggy and non-buggy methods. As a result of this enhanced separability, we believe that our approach will support improved performance in future machine learning models for bug prediction.
Summary of RQ4: After constructing the Less-Noisy dataset using the LLM-based approach, statistical analysis reveals significantly improved separability and larger effect sizes in code metrics between buggy and non-buggy methods. This suggests that the noise reduction could enhance the performance of future ML-based bug prediction models.
@@ 5. Threats to Validity
Internal Validity. The dataset used in this study was curated from method histories of Java projects obtained through CodeShovel (Grund et al., 2021), which is not fully accurate. Any inaccuracies in the method histories could impact the results. Additionally, the embedding-based approach in RQ3 relied solely on Gemini’s proprietary embedding model. Exploring embeddings generated by a broader range of LLMs may lead to more generalizable conclusions.
Conclusion Validity. All models were evaluated using standard classification metrics such as F1-score and MCC, as well as both held-out test sets and Leave-One-Out cross-validation. The conclusions drawn from these analyses depend on the quality and representativeness of the dataset. Dataset-specific characteristics may influence the observed performance trends.
@@ 6. Conclusion and Future Work
Method-level bug prediction is considered one of the holy grails in software engineering research. However, it remains an open challenge, primarily due to the absence of a noise-free method-level bug dataset (Chowdhury et al., 2024; Pascarella et al., 2020). To address this, we evaluated the effectiveness of various LLMs and prompting strategies for detecting tangled changes at the method level. Our results show that even zero-shot LLMs achieve high accuracy when both commit messages and code diffs are provided (RQ1). Performance improves further when combining chain-of-thought with few-shot prompting, surpassing the results of either technique used in isolation (RQ2). Additionally, embedding-based models deliver even greater accuracy gains (RQ3). Building on these insights, we generated a Less-Noisy dataset using our LLM-based method. This dataset shows promise in developing more accurate method-level bug prediction models as it exhibits a significantly stronger power of commonly used code metrics to differentiate between bug-prone and non-bug-prone methods compared to the original Noisy dataset (RQ4).
The implications of our findings extend beyond the untangling problem and contribute to broader discussions on the capabilities and limitations of LLMs in software engineering. Although challenges such as hallucinations, reproducibility, and data leakage remain (Liu et al., 2024b; Yao et al., 2024; Ji et al., 2023; Ye et al., 2023; Sallou et al., 2024), our study offers empirical evidence supporting the effectiveness of LLMs in tasks such as code reasoning, classification, and dataset construction (Sghaier et al., 2025; Jaoua et al., 2025; Nascimento et al., 2024). In particular, our work shows that when guided by carefully designed prompts, LLMs can significantly improve software engineering workflows.
Looking ahead, a promising direction for future research is to enrich the dataset with additional code metrics and embeddings and to explore a range of machine learning models for building more effective bug prediction models. To support continued progress, we have publicly released our code and dataset, allowing researchers to use them directly or adapt them to their specific needs.

### v3

@@ 4.3. RQ3: Performance of embedding-based machine learning models
This research question explores the viability of embedding-based ML models for detecting tangled code changes. Unlike prompt-based approaches that rely on in-context reasoning of LLMs, this method uses embeddings, vector representations of the input data from commit messages and code diffs, to train supervised classifiers. The key objective is to assess whether these embeddings, when used as features in ML models, can achieve competitive performance in detecting bug-related code changes. This experiment is motivated by the previous studies where embedding-based models performed well and even outperformed prompt-based techniques (petukhova_text_2024; Keraghel et al., 2024).
@@
Figure 3. The methodology for RQ3 has two main parts: (i) embedding generations and (ii) building ML models.:
To generate embeddings, Gemini’s text-embedding-004 was employed with the classification task type. Due to resource constraints, this study focuses exclusively on this proprietary model, utilizing Gemini’s free quota. For the open-source models, a sliding-window approach with mean pooling was applied to address token input limits, 2048 for EmbeddingGemma-300m and 512 for CodeBERT and GraphCodeBERT. All models produced 768-dimensional vector representations for each instance, derived from the concatenation of the code diff and its corresponding commit message.
The classification task was treated as a binary problem, and the dataset was split into an 80% training set and a 20% test set. We evaluated multiple classifiers, including RandomForest (RF), Support Vector Classifier (SVC), DecisionTree (DT), and Multi-Layer Perceptron (MLP). These models were selected to represent a diverse range of learning paradigms: tree-based ensembles (RF, DT), margin-based optimization (SVC), and neural networks (MLP), which have been widely applied in various software engineering tasks (pascarella_performance_2020; Chowdhury et al., 2024).
@@
Table 4. Performance comparison of ML models trained on text-embedding-004 embeddings. Metrics are calculated by targeting the buggy samples (pos_label = ”Buggy”).
Classifier Accuracy Precision Recall F1-score MCC
RF 0.816 0.841 0.82 0.83 0.629
SVC 0.884 0.918 0.866 0.891 0.768
DT 0.686 0.717 0.706 0.712 0.366
MLP 0.895 0.929 0.876 0.902 0.791
Table 4 presents the performance comparison of all models trained on text-embedding-004 embeddings, showing that the MLPClassifier achieved the best results, with an accuracy of 0.895, F1-score of 0.902, and MCC of 0.791, outperforming other classifiers in all metrics. Similar experiments were conducted using open-source embeddings; however, due to space limitations, only the MLP results are reported in Table 5, as it consistently outperformed other classifiers. Among the open-source models, GraphCodeBERT combined with MLP achieved the highest performance (F1-score = 0.831), exhibiting a modest gap of around 7% compared to the proprietary model. This suggests that while proprietary embeddings retain a advantage, open-source alternatives offer competitive and cost-effective performance for this task.
Table 5. Performance comparison of MLP classifier trained on different LLM embeddings. The first row of text-embedding-004 is taken from Table 4 for comparison. Metrics are calculated by targeting the buggy samples (pos_label = ”Buggy”).
Model Accuracy Precision Recall F1-score MCC
text-embedding-004 0.895 0.929 0.876 0.902 0.791
EmbeddingGemma 0.80 0.815 0.82 0.818 0.593
CodeBERT 0.765 0.738 0.833 0.782 0.533
GraphCodeBERT 0.828 0.831 0.83 0.831 0.656
@@
Furthermore, we evaluated the MLPClassifier using the text-embedding-004 embeddings under a Leave-One-Out (LOO) strategy to test its robustness, where a single example is used as the test sample and the remaining serve as the training data. This approach not only offers more robust evaluation, but also can show if a model’s accuracy can be improved with more training samples. Encouragingly, the model achieved improved performance with an accuracy of 0.9036, precision of 0.8980, recall of 0.9141, F1-score of 0.906, and MCC of 0.8073, setting a new benchmark compared to prior best results in RQ1 and RQ2. In terms of class-wise performance the model maintained a balanced detection capability, with precision and recall exceeding 0.89 for both classes and achieving F1-scores of 0.91 for Buggy and 0.9 for NotBuggy.
@@
Summary of RQ3: Embedding-based ML models, particularly MLPClassifier, demonstrate strong predictive power for tangled change detection, achieving over 90% F1-score and outperforming prior baselines of RQ1 and RQ2. More encouragingly, our results suggest that the performance of embedding-based models can be improved with an enlarged training dataset.
@@ 4.4. RQ4: Potential impact of LLM-based untangling on future method-level bug prediction models
From the results of RQ1 to RQ3, we observe that the LLM-based method-level tangled change detection approach performs well with high accuracy and F1-scores using both prompting techniques and embedding-based ML models. This suggests that such approaches can effectively reduce noise in bug datasets. However, an important question remains: Does this noise reduction benefit machine learning-based bug prediction models? Building such a model, however, requires addressing several challenges. First, a broad set of code metrics is needed, whereas the dataset from Chowdhury et al. (Chowdhury et al., 2024) includes only five. Second, Mashhadi et al. (Mashhadi et al., 2023) demonstrate that combining code metrics with embeddings leads to better performance, requiring us to generate embeddings for all methods. Finally, a thorough evaluation across multiple algorithms and fine-tuning strategies is essential. Due to these requirements, constructing a full bug prediction model is beyond the scope of this paper. Nonetheless, if we can demonstrate that our noise reduction approach improves the ability of code metrics to distinguish between bug-prone and non-bug-prone methods, likely, the performance of machine learning models relying on these metrics will also improve. Therefore, this research question examines whether the distribution of code metrics becomes more distinguishable in the dataset produced by our method.
In RQ1 and RQ2, we evaluated multiple prompting strategies and LLMs, showing that GPT-4o combined with few-shot + chain-of-thought prompting delivers the best and balanced performance comparing all metrics. In RQ3, we examined the feasibility of using LLM-generated embeddings with downstream machine learning classifiers to detect tangled changes. Although the latter approach yields better results, it requires generating embeddings for all tangled code diffs from bug-fix commits to support prediction with the machine learning model, which is more complex and time-consuming than the few-shot + chain-of-thought approach. Therefore, we employed GPT-4o with few-shot + chain-of-thought prompting to detect tangled changes across all methods in 49 software projects in a fast and efficient manner.
In the dataset by Chowdhury et al. (Chowdhury et al., 2024), each method is associated with its change history and the number of other methods modified in the same commit. We apply the few-shot and chain-of-thought strategy using GPT-4o to separate NotBuggy changes within bug-fix commits, thereby removing noise. To enable replication, the algorithm for creating the Less-Noisy dataset is included in the shared repository (algorithm/algorithm.png). After constructing the Less-Noisy dataset, we obtain four types of methods:
@@
The dataset provided by Chowdhury et al. (Chowdhury et al., 2024) contains five code metrics computed for each version of a method across its history. For our analysis, we consider the metrics from the first version of each method to ensure consistency across the dataset. These code metrics are defined as follows: Size refers to the number of source lines of code, excluding comments and blank lines (Chowdhury et al., 2022, 2024); Readability is a score that reflects the ease of reading the source code, developed by Buse et al. (Buse and Weimer, 2009); McCabe measures cyclomatic complexity by counting the number of independent execution paths in a method (McCabe, 1976); FanOut represents the number of distinct methods called by a given method, indicating its dependency footprint (pascarella_performance_2020; Chowdhury et al., 2024); and Maintainability Index (MI) is a composite metric that combines several aspects, including complexity and size, to offer a comprehensive assessment of code maintainability (Oman and Hagemeister, 1992).
Figure 4. Distribution of code metrics for Buggy and NotBuggy methods in the Titan project across the Noisy and Less-Noisy datasets. Metric distributions differ more significantly in the Less-Noisy dataset.:
@@
To further quantify the differences between Buggy and NotBuggy methods, we compute Cliff’s Delta effect sizes across all 49 projects. Table 6 presents the effect size distributions for both the Noisy and Less-Noisy datasets. We categorize the effect sizes following Hess et al. (Hess and Kromrey, 2004): Negligible (<0.147<0.147), Small ($0.147\leq\delta<0.33$), Medium ($0.33\leq\delta<0.474$), and Large ($\delta\geq 0.474$).
@@
The table reports the proportion of projects falling into each effect size category for all five code metrics. The results indicate that, for most metrics, the Less-Noisy dataset shows a substantial shift toward larger effect sizes. For example, for the code metric, Size, in the Less-Noisy dataset, 73% of the projects exhibit a large effect size in the difference between Buggy and NotBuggy methods, compared to only 36% in the Noisy dataset. Similar patterns are observed for the other four code metrics. These findings clearly demonstrate that our filtering approach effectively eliminates noise and increases the distinction between the code metrics of buggy and non-buggy methods. As a result of this enhanced separability, we believe that our approach will support improved performance in future machine learning models for bug prediction.
Summary of RQ4: After constructing the Less-Noisy dataset using the LLM-based approach, statistical analysis reveals significantly improved separability (with larger effect sizes) in code metrics between buggy and non-buggy methods. This suggests that the noise reduction could enhance the performance of future ML-based bug prediction models.
@@ 5. Threats to Validity
Internal Validity. The dataset used in this study was curated from method histories of Java projects obtained through CodeShovel (Grund et al., 2021), which is not fully accurate. Any inaccuracies in the method histories could impact the results.
Conclusion Validity. It is affected by all of the above mentioned threats.
@@ 6. Conclusion and Future Work
Method-level bug prediction is considered one of the holy grails in software engineering research. However, it remains an open challenge, primarily due to the absence of a noise-free method-level bug dataset (Chowdhury et al., 2024; pascarella_performance_2020). To address this, we evaluated the effectiveness of various LLMs and prompting strategies for detecting tangled changes at the method level. Our results show that even zero-shot LLMs achieve high accuracy when both commit messages and code diffs are provided (RQ1). Performance improves further when combining chain-of-thought with few-shot prompting (RQ2). Additionally, embedding-based models deliver even greater accuracy gains (RQ3). Across RQ1 to RQ3, open-source models also demonstrate strong performance. While they do not yet match proprietary models, they offer efficient alternatives with acceptable trade-offs. Building on these insights, we generated a Less-Noisy dataset using our LLM-based method. This dataset shows promise in developing more accurate method-level bug prediction models as it exhibits a significantly stronger power of commonly used code metrics to differentiate between bug-prone and non-bug-prone methods compared to the original Noisy dataset (RQ4).
The implications of our findings extend beyond the untangling problem and contribute to broader discussions on the capabilities and limitations of LLMs in software engineering. Although challenges such as hallucinations, reproducibility, and data leakage remain (yao_llm_2024; Ji et al., 2023; ye_cognitive_2023; sallou_breaking_2024), our study offers empirical evidence supporting the effectiveness of LLMs in tasks such as code reasoning, classification, and dataset construction (zheng2024few; Jaoua et al., 2025; Nam et al., 2024). In particular, our work shows that when guided by carefully designed prompts, LLMs can significantly improve software engineering workflows.
Looking ahead, a promising direction for future research is to enrich the dataset with additional code metrics and embeddings and to explore a range of machine learning models for building more effective method-level bug prediction models. To support continued progress, we have publicly released our code and dataset, allowing researchers to use them directly or adapt them to their specific needs.